### v1

# Rapid Parameter Estimation for Merging Massive Black Hole Binaries Using ODE-Based Generative Models

Abstract
Detecting the coalescences of massive black hole binaries (MBHBs) is one of the primary targets for space-based gravitational wave observatories such as LISA, Taiji, and Tianqin. The fast and accurate parameter estimation of merging MBHBs is of great significance for both astrophysics and the global fitting of all resolvable sources. However, such analyses entail significant computational costs. To address these challenges, inspired by the latest progress in generative models, we proposed a novel artificial intelligence (AI) based parameter estimation method called Variance Preserving Flow Matching Posterior Estimation (VPFMPE). Specifically, we utilize triangular interpolation to maintain variance over time, thereby constructing a transport path for training continuous normalization flows. Compared to the simple linear interpolation method used in flow matching to construct the optimal transport path, our approach better captures continuous temporal variations, making it more suitable for the parameter estimation of MBHBs. We also constructed a novel network architecture for gravitational wave feature extraction that outperforms existing methods in terms of training speed and accuracy. Additionally, we creatively introduce a parameter transformation method based on the symmetry in the detector’s response function. This transformation is integrated within VPFMPE, allowing us to train the model using a simplified dataset, and then perform parameter estimation on more general data, hence also acting as a crucial factor in improving the training speed. In conclusion, for the first time, within a comprehensive and reasonable parameter range, we have achieved a complete and unbiased 11-dimensional rapid inference for MBHBs in the presence of astrophysical confusion noise using ODE-based generative models. In the experiments based on simulated data, our model produces posterior distributions comparable to those obtained by nested sampling.
1 Introduction
Gravitational wave (GW) astronomy has made significant progress due to the breakthroughs in ground-based detections led by the LIGO-Virgo-KAGRA network [1, 2, 3]. Scheduled in the upcoming decade, space-based detectors such as the Laser Interferometer Space Antenna (LISA) [4, 5], Taiji [6, 7, 8], and Tianqin [9, 10] aim to detect GWs in the 0.1 mHz - 1 Hz frequency band associated with enormous astrophysical and cosmological sources [11]. As one of the representative projects, Taiji consists of a triangle of three spacecraft (S/Cs) with a baseline separation of 3 million kilometers and aims to detect low-frequency GWs emitted by sources such as compact galactic binaries (GBs) [12], massive black hole binaries (MBHBs) [13], extreme mass ratio inspirals (EMRIs) [14], as well as the stochastic gravitational wave background of astrophysical or cosmological origins.
The merger of MBHB with component masses $10^{4}\sim 10^{7}M_{\odot}$ is one of the primary observation targets of space-based GW detectors, since they may shed light upon the growth and merger history of massive black holes, the dynamic behavior of curved spacetime, and the nature of gravity, etc. Due to the presence of the surrounding gas environment, MBHBs may produce detectable electromagnetic (EM) signals in addition to GWs before, during, and after the mergers [15, 16, 17, 18]. Joint EM and GW observation of the same source may unveil the mystery of the evolution of MBHBs and their host galaxies, and provide a new measurement of the cosmic expansion rate, hence shedding light upon the Hubble tension problem in cosmology [19, 20]. For these purposes, it is crucial that the estimation of MBHB parameters, especially the sky location (ecliptic longitude, latitude), should be conducted with high precision and low latency, to guide the search of EM observatories.
Besides, the parameter estimation of merging MBHB must also be robust against the contamination of the confusion foreground formed by tens of millions of Galactic binaries. Indeed, in the area of space-based GW detection, data processing faces the challenge caused by the presence of a large number of astrophysical sources in the measurement band [21]. The non-trivial overlap of individual signals in the time and frequency domains increases the complexity of the task. The accurate analysis of up to $\mathcal{O}(10^{4})$ overlapping sources in the data requires the use of global fit techniques [22, 23, 24], namely jointly fitting all sources and noise in the data, which imposes a high challenge on both computational time and resources. The prototype global fit pipeline proposed by Ref. [24] currently requires as many as $\mathcal{O}(10^{3})$ CPUs and several days to process the 12-month Lisa Data Challenge (LDC) data, and the issue of computational resources would become even more critical for the multi-year datasets in the future. Due to these challenges and problems, the aim of this paper is to develop an accelerated and robust method to estimate the parameters of merging MBHBs.
When discussing the efficiency and speed of parameter estimation, the advancement in machine learning technology is revolutionizing our understanding of this area. Currently, most machine learning methods applied to the merging of MBHB use neural posterior estimation(NPE) [25, 26, 27] of discrete normalized flows. NPE has been widely used in fields such as atmospheric retrievals (AR) [28] and ground-based gravitational wave parameters estimation [29, 30]. NPE was also used by the winning entry to the 2023 edition of the ARIEL data challenge [31].
The mission lifetimes of LISA, Taiji, and Tianqin are planned to last for years, during which the payloads, orbits, and noise characteristics will vary over time. Accordingly, the models used for parameter estimation may also need to be regularly updated. As a result, the training speed of the model will be linked to the timeliness of scientific output. However, a significant challenge to the current NPE model in MBHB inference is the extensive training duration, which hampers its ability to conduct real-time data analysis. Additionally, previous research [32] in this area shows that NPE falls short in accurately reproducing the posterior distribution obtained by the benchmark Nested Sampling method. This prolonged training period not only delays the application of the model but also limits its scientific utility in rapidly evolving research contexts. To address these challenges, we applied the latest generative model to MBHBs and made further improvements and optimizations to better solve the problem. Compared to NPE, Flow Matching Posterior Estimation (FMPE) offers more flexibility in designing non-diffusion paths, such as optimal transport paths, and enables direct access to density functions [33]. Moreover, this method can achieve faster and more accurate results than NPE. Nevertheless, FMPE relies solely on simple linear interpolants when constructing optimal transport paths. This reliance impacts the extraction and modeling of MBHB waveform features, which in turn affects the efficiency and accuracy of parameter estimation.
Therefore, to capture the continuous temporal variations and enhance the parameter estimation of MBHBs, we aim to explore more effective methods for constructing transport paths and MBHB wave source feature extraction, which are particularly beneficial and targeted for extracting the continuous waveform characteristics of MBHBs. We consider utilizing trig interpolants to design transport paths [34, 35] that allow for the continuous evolution of MBHB waveform features along the time axis for modeling. This modeling accurately bridges any two probability density functions within a finite time, which belongs to a broad category of continuous-time stochastic processes. We call this method Variance Preserving Flow Matching Posterior Estimation (VPFMPE). Compared to standard Monte Carlo methods, the VPFMPE model achieves comparable performance.
Figure 1: The Diagrams for the training and sampling phases of our model. On the lower left, the training phase is displayed, where $t$ denotes time and $x_{1}$ represents the data from the first day of observed MBHBs. In this phase, a feature extractor, as detailed in section 3.1, processes the data to produce a final vector field. This vector field enables the network architecture to learn a transformation from a base distribution ($\theta_{0}$, sampled from a Gaussian distribution) to a posterior distribution ($\theta_{1}$, sampled from the posterior distribution). On the lower right, the sampling phase is shown. Here, the data with noise ($x_{t}$), which includes both confusion noise and instrumental noise from MBHB observations at any time, is mapped back to the first day’s data $x_{1}$ using the coordinate transformation described in section 3.2. The Gaussian distribution is then converted into a posterior distribution by solving the ordinary differential equation (ODE) with the vector field learned during the training phase. The top of the image displays simulation data from both the training and sampling phases.
The challenges of parameter estimation for MBHBs not only involve the need for sophisticated algorithms capable of advanced feature extraction and inference, but also lie in the excessively wide prior range of parameters and the resulting complexity of data. Recent studies [36, 32] adopted priors that do not fully cover the complete physical ranges of parameters. This compromise was due to that setting broad prior ranges during model training can reduce the effectiveness of the model. This underscores the importance of finding a careful balance between the breadth of prior ranges and the accuracy of estimation. To date, no research has managed to achieve unbiased parameter estimation across the entire range of physical priors. Our method is proven to overcome this limitation. In this paper, we expanded the prior ranges of parameters to improve the applicability model.
In order to further enhance the efficiency and effectiveness of model training, we tackled a common complexity faced by space-based gravitational wave detectors, namely the time-varying response function. Utilizing the coordinate transformation of parameters, we can essentially conduct both the model training and the parameter estimation in a coordinate frame fixed to the detector, in which the response function is time-invariant. This strategy significantly simplifies the training task. Consequently, the integration of coordinate transformation and advanced generative models enhances the speed and accuracy of parameter estimation for MBHBs across an extensive range of priors.
In addition, we have developed a novel network architecture for gravitational wave feature extraction. This architecture illustrated in Figure 4, named TResMlp, introduces the innovation of encoding time $t$, resulting in superior training speed and accuracy compared to existing methods [33, 32, 36].
The training of the model can be finished in about two days, as shown in Figure 1, which is roughly three times faster than the training duration in our previous work [32]. Additionally, its performance on space-based gravitational wave data is on par with that of traditional Monte Carlo methods. This demonstrates the significant potential of integrating artificial intelligence with VPFMPE techniques to achieve results that are comparable to those obtained with standard Monte Carlo methods.
2 Data Generation and Preprocessing
In this section, we briefly describe the simulation of data used in model training and validation. The IMRPhenomD model is adopted to describe the source-frame waveforms of MBHBs. Given the widely recognized necessity of using Time Delay Interferometry (TDI) to suppress laser frequency noise, Taiji’s response to the MBHB signals is calculated in the form of two noise-orthogonal TDI channels $(A,E)$. Specifically, the fast Fourier domain transfer function derived by Ref. [37] is employed to project the IMRPhenomD waveform to the $(A,E)$ channels, which can be written in a concise form as
$\tilde{A}(f)=\sum_{\alpha}\mathcal{T}^{A}_{\alpha}(f,t_{f})\tilde{h}_{\alpha}(f),\quad\tilde{E}(f)=\sum_{\alpha}\mathcal{T}^{E}_{\alpha}(f,t_{f})\tilde{h}_{\alpha}(f),\quad\alpha\in\{+,\times\}.$ (1)
As can be seen, the transfer function has a close dependence on the orbits of the detector (through the $t_{f}$ term). For the sake of speed and simplicity, the motions of Taiji’s satellites are modeled based on an equal-arm analytic orbit with nominal arm-length $L=3\times 10^{9}{\rm m}$ and an orbital period of 1 year.
The TDI response of each MBHB signal can be fully specified by 11 parameters: the chirp mass $\mathcal{M}_{c}$, mass ratio $q\equiv m_{2}/m_{1}$, the dimensionless spins along the $z-$axis $\chi_{z1}$ and $\chi_{z2}$, the time $t_{c}$ and phase $\varphi_{c}$ at coalescence, the luminosity distance $d_{L}$, the inclination angle $\iota$, polarization angle $\psi$, and the ecliptic coordinate $\{\lambda,\beta\}$, being the longitude and latitude respectively. The priors on these parameters are listed in Table 1. Similar to the LDC-sangria data, the prior range of $\mathcal{M}_{c}$ covers an order of magnitude. In our previous work [32], considering the complexity of data, we limited the prior ranges of spins to positive values. On this simplified parameter estimation task, NPE has been demonstrated to achieve unbiased estimations. However, its performance is subpar when applied to MBHB data with a complete spin prior. An advantage of the models proposed in this paper is that they are applicable to the complete physical prior range of spins (i.e. $[-1,1]$). Parameter $t_{c}$ is defined relative to a reference time $t_{\rm ref}$. Since the mergers of MBHBs manifest as prominent peaks in the time domain, we consider a prior range of $\sim 2000$ s ( $\sim 0.02$ day) around $t_{\rm ref}$. Notably, to simplify the dataset and reduce the difficulty of training, $t_{\rm ref}$ is fixed to the $1^{\rm th}$ day in the training set. The response function varies with the motion of detectors with a 1-year period, thus a training set with $t_{c}\in 1\pm 0.01$ day is far from being sufficient to cover all the observable MBHBs. While, thanks to the method described in Sec. 3.2, our model can be generalized to make inferences throughout the entire task cycle. Besides, the range of $d_{L}$ corresponds to the redshift range of $z\in[1,10]$, which is sufficient to encompass the majority of MBHBs observable by space-based GW detectors. The angles $\iota$ and $\beta$ are uniformly distributed in terms of $\cos\iota$ and $\sin\beta$ to represent that the sources are isotropically distributed on the sphere.
Table 1: The prior ranges of parameters. The last column is the alias of parameters used in the code as well as the corner plot.
Parameter Description Prior Lower Bound Prior Upper Bound Alias in Code $\mathcal{M}_{c}$ Chirp mass of the binary system $1.0\times 10^{5}M_{\odot}$ $1.0\times 10^{6}M_{\odot}$ chirp_mass $q$ Ratio of masses of the two black holes 0.1 1.0 mass_ratio $\chi_{z1}$ Spin of the 1st black hole along the $z$-axis -1 1 spin_1 $\chi_{z2}$ Spin of the 2nd black hole along the $z$-axis -1 1 spin_2 $t_{c}$0.01 Time of coalescence relative to the reference time  [^1] The reference time $t_{\rm ref}$ is set to the $1^{\rm th}$ day., $t_{\rm ref}$ day 0.01 day coalescence_time $\varphi_{c}$ Phase at the moment of coalescence 0 rad $2\pi$ rad coalescence_phase $d_{L}$ Luminosity distance to the binary system 6000 Mpc 100000 Mpc luminosity_distance $\iota$ Angle of inclination of the binary orbit 0 rad $\pi$ rad inclination $\beta$ Ecliptic latitude of the binary system $-\frac{\pi}{2}$ rad $\frac{\pi}{2}$ rad latitude $\lambda$ Ecliptic longitude of the binary system 0 rad $2\pi$ rad longitude $\psi$ Polarization angle of the gravitational wave 0 rad $\pi$ rad psi
The majority of MBHB’s signal-to-noise ratio(SNR) is contributed by a short duration including the merger phase (typically several hours) [38]. To reasonably avoid the potential overlap of other signals, the time duration for each sample is set to one day, with the merger positioned towards the end of this window. Furthermore, the sampling frequency is set as $1/15$ Hz, indicating a Nyquist frequency of $0.033$ Hz. This frequency exceeds the highest instantaneous frequency of the waveforms, ensuring comprehensive capture of the signal.
Figure 2: A representative data in the test set. The orange and green curves represent the Fourier transforms of the TDI-A, E data streams, which include GW signals, instrumental noises, and confusion noise. To clearly illustrate the relative strengths of the signal and noises, the blue curve depicts the Fourier transform of the confusion noise, while the gray dashed curve indicates the theoretical amplitude of the instrumental noises.
Along with the detector responses, we also need the PSD of the noise in each channel. For instrumental noises, according to the current design of the Taiji mission, the one-sided ASDs of the two top-level instrumental noises, namely optical measurement system noise and test mass acceleration noise, are  [39]
$\sqrt{S_{\mathrm{OMS}}(f)}=$ $8\times 10^{-12}\sqrt{1+\left(\frac{2\mathrm{mHz}}{f}\right)^{4}}\frac{{\mathrm{m}}}{\sqrt{\mathrm{Hz}}},$ (2)
$\sqrt{S_{\mathrm{ACC}}(f)}=$ $3\times 10^{-15}\sqrt{1+\left(\frac{0.4\mathrm{mHz}}{f}\right)^{2}}\sqrt{1+\left(\frac{f}{8\mathrm{mHz}}\right)^{4}}\frac{{\mathrm{m/s^{2}}}}{\sqrt{\mathrm{Hz}}},$ (3)
The total instrumental noise PSDs of the $A,E$ channels $S_{n}^{A,E}(f)$ are the combinations of these two components (see Ref. [40, 41, 42] for their specific forms). For simplicity, we treat $S_{n}(f)$ as constant over the whole mission duration. Our training data is a combination of “clean” TDI responses and stationary instrumental noises. On a more realistic ground, the long-term variations and short-term jitters of instrumental noise, along with the typical anomalies such as glitches and gaps should be considered in future work.
To test the robustness and generalizability of the model, during the inference stage, we consider a non-stationary noise caused by the unresolvable foreground originating from the overlapping GWs of GBs. This confusion noise is generated with the FastGB algorithm [43] from a catalog of $\sim 3\times 10^{7}$ GBs provided by the LDC dataset, from which $\sim 2\times 10^{4}$ bright sources with SNR greater than 7 were subtracted. This choice of SNR threshold follows the convention of [44, 45, 46, 47]. Shown in Figure 2 is representative data in the test set, including GW signal, instrumental noises, and confusion noise.
3 Methodological Framework
The application of machine learning techniques in the field of gravitational waves has seen widespread success and remarkable achievements [48, 49]. Among these techniques, deep learning has demonstrated exceptional prowess, especially in processing ground-based gravitational wave data [50, 51, 52, 53, 54, 55, 56, 57, 58].
NPE with normalizing flows. Neural Posterior Estimation (NPE) combined with Normalizing Flows (NF) is used to directly fit the posterior distribution $p(\theta|x)$. In this approach, an invertible mapping $\psi_{x}$ transforms the base distribution $p_{0}(\mathbf{z})$ into the target distribution $q(\theta|x)$:
$q(\theta|x)=p_{0}(\psi_{x}^{-1}(\theta))\det\left|\left(\frac{\partial\psi_{x}^{-1}(\theta)}{\partial\theta}\right)\right|.$ (4)
NPE is trained using a maximum likelihood objective and simplifies the expectation using Bayes’ theorem.
$L_{\mathrm{NPE}}=-E_{p(\theta)p(x|\theta)}\log q(\theta|x).$ (5)
Once training is complete, NPE can efficiently perform inference on new observations $x$ and provide exact density evaluations of $q(\theta|x)$. This method retains the expressive power and computational efficiency of normalizing flows while achieving posterior fitting.
Continuous Normalizing Flows for parameter estimation. Continuous Normalizing Flows (CNFs) transform a simple base distribution into a more complex distribution by describing this transformation through continuous time $t\in[0,1]$. For each $t$, the flow is defined by a vector field $v_{t,x}$ on the sample space. This vector field represents the trajectory velocity of the samples at any given time $t$.
$\frac{d}{dt}\psi_{t,x}(\theta)=v_{t,x}(\psi_{t,x}(\theta)).$ (6)
The conversion between the base distribution (at $t=0$) and the target distribution (at $t=1$) is achieved by integration. A straightforward approach to training Continuous Normalizing Flows (CNFs) involves starting with a given base distribution and using an ODE solver to obtain the target distribution. This target distribution is then constrained to match the real data distribution by minimizing a divergence measure, such as the KL divergence.
Because many intermediate trajectories are unknown, inferring the distribution (either through sampling or likelihood computation) requires repeatedly simulating the ODE, leading to a significant computational burden.
However, Flow matching provides alternative training objectives for CNFs, Dax et al. [33] apply flow matching [59] to simulation-based inference. The loss function of FMPE is expressed as:
$L_{FMPE}=E_{t\sim p(t),\theta_{1}\sim p(\theta),x\sim p(x|\theta_{1}),\theta_{t}\sim p_{t}(\theta_{t}|\theta_{1})}\parallel v_{t,x}(\theta_{t})-u_{t}(\theta_{t}|\theta_{1})\parallel^{2}.$ (7)
Here, $v_{t,x}(\theta_{t})$ represents the vector field that generates the path to the target probability distribution, and $u_{t}(\theta_{t}|\theta_{1})$ is the sample-conditioned vector field. FMPE introduces a Gaussian path with time-dependent means $t\theta_{1}$ and standard deviations $1-(1-\sigma)t$.
$p_{t}(\theta|\theta_{1})=N(t\theta_{1},1-(1-\sigma)t).$ (8)
FMPE defines the sample-conditioned vector field is
$u_{t}(\theta|\theta_{1})=\frac{\theta_{1}-(1-\sigma)\theta_{0}}{1-(1-\sigma)t}.$ (9)
In our work, we reference the studies by Albergo et al. [34, 35] and utilize triangular interpolation to define the time-dependent mean and standard deviation, thus constructing the Gaussian path and sample-conditioned vector field as follows:
$p_{t}(\theta|\theta_{1})=N(\sin(\frac{\pi}{2t})\theta_{1}+\cos(\frac{\pi}{2t})\theta_{0},\sigma).$ (10)
$u_{t}(\theta|\theta_{1})=\frac{\pi}{2(\cos(\frac{\pi}{2t})\theta_{1}-\sin(\frac{\pi}{2t})\theta_{0})}.$ (11)
We call this method Variance Preserving Flow Matching Posterior Estimation and use it to train continuous normalization flows. We show the differences between the two methods in Table  2.
Table 2: A table with constructing transport paths differences. FMPE uses linear interpolation to construct transport paths, while VPFMPE uses trig interpolants to construct transport paths.
Interpolar Method Mean of the probability path Standard deviation of the probability path
FMPE linear $t\theta_{1}$ $1-(1-\sigma)t$
VPFMPE trig $\sin(\frac{\pi}{2t})\theta_{1}+\cos(\frac{\pi}{2t})\theta_{0}$ $\sigma$
Figure 3: Comparison of the posterior distributions of the two methods. Blue represents the posterior distribution of the VPFMPE method on the injected MBHB signal. Orange represents the posterior distribution of the FMPE method on the injected MBHB signal. The red line represents the true value of the injected MBHB signal.
Comparison of methods. To demonstrate the performance of the transport path constructed by VPFMPE, we randomly injected an MBHB signal and compared the FMPE method with the VPFMPE method. Both models were trained using the same learning rate, optimizer, and gravitational wave feature extractor to ensure fairness. We show in Figure 3 a comparison of the posterior estimates of the four parameters of the injected MBHB signal by the two methods.
Discussion. The NPE network just only one-third allocated to the feature extractor network for GW data. However, the ODE-based generative model only needs to parameterize a simple vector field, so it can devote its full network capacity to interpreting the higher dimensions [33].
3.1 Network architecture
A key challenge in space-based gravitational wave parameter estimation lies in managing the high-dimensional nature of its data. To address this, it is essential to develop a feature extraction network capable of transforming the high-dimensional input data into a more manageable, low-dimensional representation [29, 32, 60]. This transformation is critical for adjusting the dimensionality of the final feature vector to match that of the parameter vector $\theta$, enabling the network to effectively parameterize the conditional vector field $(t,x,\theta)\rightarrow v_{t,x}(\theta)$ [33]. Achieving accurate parameterization of this conditional vector field is fundamental to the construction of an effective network structure.
Figure 4: TResMlp Network Architecture: Initially, the frequency vector $f$ is used to encode the position of time $t$. Sine and cosine functions are then applied to extract periodic features from time $t$, resulting in the “Time Embedding”. Subsequently, a “Linear” layer maps the two channels of the space-based GW data to produce the “Data Embedding”. Finally, all obtained embeddings are input into a Multi-Layer Perceptron (MLP) with a residual structure to generate the final output $v_{(t,x_{1})}(\theta_{t})$.
In recent years, DeResNet networks have been commonly used as the backbone network for feature extraction in ground-based gravitational wave parameter estimations or space-based gravitational wave parameter estimation [57, 32, 33, 53], and Gated Linear Unit(GLU)-conditioning [61] can be optionally used for information aggregation of $(\theta,t)$.
However, if information about the actual parameter $\theta$ is not properly handled or retained during the process of aggregating $(\theta,t)$ information using GLU, it could potentially lead to a loss of information about $\theta$. To encode the time $t$ more effectively and ensure that our feature extraction network captures its dynamic nature, we introduce a frequency vector $f$. This frequency vector $f$ adapts dynamically to changes in the parameters of the wave source, allowing for a more precise encoding of time $t$. To capture the periodic features of time, we employ sine and cosine functions for encoding. Our network architecture, referred to as TResMlp, is shown in Figure 4.
We trained the CNFs of MBHBs using two different network architectures. To ensure fairness, the two models had identical parameters such as the optimizer and learning rate, differing only in their network structures. Additionally, both architectures used the same hidden layer configuration. We show in Figure 5 a comparison of the posterior estimates of the four parameters of the injected MBHBs signal by the two models.
Figure 5: Comparison of the posterior distributions of the two network architectures. The blue represents the posterior distribution plot of CNFs trained using the TResMlp network. The orange represents the posterior distribution plot of CNFs trained using the DeResNet network. The red line represents the true value of the MBHBs.
Compared to the DeResNet model that use GLU-conditioning for aggregating information about $(\theta,t)$ [57, 32, 33, 53], our network architecture, TResMlp, demonstrates faster training speed and higher precision with the same hidden layers. More experiments and comparisons of related models are described in detail in Section 4.
3.2 Coordinate Transformation and Model Generalization Strategy
To simplify the dataset and reduce the difficulty of training, the coalescence times of MBHBs are set within $1\pm 0.02$ day in the training set (see Sec. 2). While the response function of detector to GW signals varies in time with a one-year period, thus in principle the model trained on this dataset can only be applied to MBHBs that merge within $1\pm 0.02$ day. While, thanks to the generalization strategy in this section, the application scope of the model can be extended to the whole mission lifetime.
In this paper, the time of coalescence (dubbed $t_{c}$) is defined relative to a reference time $t_{\rm ref}$. The complexity of space-based GW data partly lies in the annual modulation of detector response due to the orbital motion of detector constellation around the sun. In other words, the mapping between GW data and MBHB parameters varies with $t_{\rm ref}$. This significantly increases the complexity of the training set, which would severely damage the training efficiency and the performance of the model. To tackle this challenge, we exploit the symmetry (invariance) of the system, fix the training data around a certain $t_{\rm ref0}$, and then the trained model is generalized via our transformation rule to make inferences at any time. The idea behind this is a paradigm that combines physical information with neural networks, and methods belonging to this family, such as the group-equivariant neural posterior estimation [62, 60], have shown broad application prospects.
Figure 6: Comparison of the effect of the model generalization strategy on model performance. The left panel shows the P-P plot without the model generalization strategy, while the right panel shows the P-P plot with the model generalization strategy. Both plots are for a set of 4000 injections with added instrument noise and reference times ranging from 1 to 365 days. Ideally, the cumulative distribution function of these percentiles for each parameter should align closely with the diagonal line, indicating optimal network performance. In such a case, the percentiles would be uniformly distributed between 0 and 1. The grey regions on the plot represent the $1\sigma$ and $2\sigma$ confidence intervals.
We describe the transformation rule as follows. We define two coordinate systems, namely the Solar System Barycenter (SSB), in which physical quantities are indexed with “S”, and the Taiji detector reference frame, in which physical quantities are indexed with “T”. The parameters which vary with the coordinate systems are $\{t_{c},\lambda,\beta,\psi\}$, where $t_{c}$ should be interpreted as the time of arrival at the origin of the coordinate system (relative to $t_{\rm ref}$). Since the data in consideration last only for one day, for preliminary parameter estimation tasks, the location of MBHB remains stationary relative to the Taiji reference frame, and the mapping between source parameters $\bm{\theta}_{T}$ and TDI response is independent of $t_{\rm ref}$. This is precisely the symmetry based on which we can achieve model generalization. The $t_{\rm ref}$-dependent coordinate transformation from Taiji frame to SSB frame is denoted as $f_{t_{\rm ref}}:\bm{\theta}_{T}\rightarrow\bm{\theta}_{S,{t_{\rm ref}}}$ (see Ref. [63] for the expression of $f$). The parameters are firstly inferred with the model trained around $t_{\rm ref0}$, hence dubbed $\bm{\theta}_{S,t_{\rm ref0}}$, and the targeted parameters $\bm{\theta}_{S,t_{\rm ref}}$ can be obtained following two steps:
$\bm{\theta}_{T}=f_{t_{\rm ref0}}^{-1}\left(\bm{\theta}_{S,t_{\rm ref0}}\right),\quad\bm{\theta}_{S,t_{\rm ref}}=f_{t_{\rm ref}}\left(\bm{\theta}_{T}\right),$ (12)
or in terms of posterior distributions:
$p\left(\bm{\theta}_{S,t_{\rm ref}}|d\right)=p\left(\bm{\theta}_{S,t_{\rm ref0}}|d\right)J\left(f^{\ }_{t_{\rm ref0}}\right)J\left(f^{-1}_{t_{\rm ref}}\right),$ (13)
with $J\left(f^{\ }_{t_{\rm ref0}}\right)$ and $J\left(f^{-1}_{t_{\rm ref}}\right)$ being the Jacobian determinants corresponding to $f_{t_{\rm ref0}}$ and $f^{-1}_{t_{\rm ref}}$, respectively. Formula (13) looks similar to the transformations in normalizing flows, except that the mapping $f_{t_{\rm ref}}$ is constructed based on our understanding of the system, instead of training a neural network. Training a network to represent this transformation is unnecessary since we possess comprehensive knowledge of its form and can effortlessly compute it. As a result, The embedding of this physical information greatly simplifies training tasks.
Below, we describe how models can be combined with the model generalization strategy. Firstly, during the dataset generation, the merger times of MBHBs are limited to 0.02 days centered at a specific reference time (the 1st day). This significantly simplifies the dataset’s complexity and reduces the challenge of MBHB parameter estimation. Subsequently, we employ VPFMPE as the loss function for training CNFs. We utilize the novel network architecture TResMlp as the gravitational wave feature extractor, labeling this approach as VPFMPE-TResMlp.
In the inference phase, we use the model generalization strategy to map the reference time data for inference to the reference time data fixed during training. This allows the model to be trained on only one day of data but tested and applied to data spanning an entire year. We use the P-P plot in Figure 6 to demonstrate the importance of using a model generalization strategy.
4 Experiments
To ensure unbiased parameter estimation, we examine our algorithm on injected GW signals and noise in this section. For the parameter estimation of the signals, we employed the four models mentioned above: VPFMPE-TResMlp, FMPE-TResMlp, FMPE-DeResNet [33], and NPE-DeResNet [32]. We applied the model generalization strategy 3.2 to all four models by default. Otherwise, due to the complexity of the dataset, the models would lack inference capability.
Table 3: In the 850 simulations, we compared the correlation coefficients between the true values and the recovered values for the four models. In addition, in the last column of the table, we have included the training time for the three models.
$\mathcal{M}_{c}$ $q$ $\chi_{z1}$ $\chi_{z2}$ $d_{L}$ Training time
VPFMPE-TResMlp 0.9978 0.9790 0.9723 0.8163 0.9408 $\approx$ 2.0 days
FMPE-TResMlp 0.9978 0.9787 0.9719 0.8139 0.9400 $\approx$ 2.0 days
FMPE-DeResNet 0.9970 0.9762 0.9563 0.7174 0.9397 $\approx$ 2.3 days
NPE-DeResNet 0.9261 0.8647 0.6620 0.1681 0.8634 $\approx$ 6.5 days
To demonstrate the effectiveness of our algorithm and validate that the recovered probability distribution accurately represents the confidence we should have in the signal parameters, we conducted extensive testing on simulated signals and noise. We injected 850 waveforms drawn from the prior, along with residual instrumental and foreground confusion noise. We will take the median of the recovered values as the point estimate and calculate the correlation coefficient with the true values of the parameters. The correlation coefficient reflects the degree of linear correlation between the true value distribution and the estimated values. A coefficient value of 1 indicates perfect consistency between the estimated and true values. Note that the reference times of these testing signals are randomly distributed throughout the year, rather than being fixed to the 1st day.
Our results are presented in Table 3. In the final column of the Table 3, we have included the approximate training times for the three models. The VPFMPE-TResMlp model achieved the best results and had the fastest training time. For $t_{c}$, $\varphi_{c}$, $\iota$, $\beta$, $\lambda$, and $\psi$, due to the presence of multimodality (i.e. there is more than one peak in the posterior distributions), using the median as a point estimate is not applicable.
Figure 7: Corner plot of parameter posterior distributions for gravitational wave signals injected with instrumental and confusion noise. The blue contours illustrate the two-dimensional joint posteriors obtained from our (VPFMPE-TResMlp) model, with peaks selected based on $t_{c}$ that align with true values. In contrast, the orange contours represent the corresponding posteriors obtained from benchmark analysis using Nested Sampling. For each model, the contour boundaries respectively encompass 95% and 60% confidence levels. The true parameter values of the simulated signal are indicated by red vertical and horizontal lines.
To date, existing research has not yet succeeded in providing unbiased estimates of MBHBs across the entire parameter space. For example, Ref. [36] only focused on four parameters, and Ref. [32] have limited to positive spins. Also, we use machine learning techniques for complete parameter estimation of MBHBs and achieve comparable results to nested sampling. Therefore, in subsequent comparisons, we will only compare the results of VPFMPE-TResMlp to those of the MCMC method.
The aim of our ODE-based generative models for MBHBs is to reproduce the comparable posterior distribution as the MCMC method. Thus, by comparing the results of Bayesian inference on the same event, the performance of the model can be better assessed. In this case, we use Nested Sampling [64, 65, 66] as a Monte Carlo technique for comparison experiments, as shown in Figure 7.
An illustration demonstrating an example using the same prior probabilities outlined in Table 1 is provided in Figure 7, where a test injection is subjected to confusion noise on the 30th reference day. The injected GW signal originates from an MBHB with a total signal-to-noise ratio of 1543, and its parameters are detailed in Table 4. Compared to the standard nested sampling technique, our model exhibited results similar to the traditional nested sampling, demonstrating consistency in the sense of the degeneracy among extrinsic parameters 4. This consistency aligns with the crucial physical characteristics of MBHBs. Note that compared to nested sampling, our machine learning process always produces similar posterior distributions on $t_{c}$, $d_{L}$, $\iota$, $\beta$, $\lambda$, and $\psi$ 7.
We have achieved accurate estimates of the sky position parameters. This is most advantageous since accurate and reliable estimates of the sky position parameters (longitude, latitude) are useful in guiding the search for electromagnetic observatories.
We have collated the true values of the injections alongside the results generated by both the model and the nested sampling process in Table  4. This provides a concise overview of the performance of each method.
Table 4: This table compares the parameters injected in the prior with the parameters recovered by the Nested Sampling and VPFMPE-TResMlp methods. The GW signal is emitted by a merging MBHB with a total SNR of 1543. The recovered values are accompanied by their 1 $\sigma$ confidence regions.
Parameter Injected value VPFMPE-TResMlp Nest
$\mathcal{M}_{c}$ [$M_{\odot}$] 531067.4 $528444.8000_{-1206.2360}^{+1141.3900}$ $531038.3974_{-107.2206}^{+125.0852}$
$q$ 0.6355 $0.5994_{-0.0220}^{+0.0236}$ $0.6405_{-0.0070}^{+0.0060}$
$\chi_{z1}$ 0.2930 $0.4051_{-0.1042}^{+0.0858}$ $0.2636_{-0.0259}^{+0.0318}$
$\chi_{z2}$ 0.2726 $0.0273_{-0.1552}^{+0.1807}$ $0.3233_{-0.0517}^{+0.0409}$
$t_{c}$ [day] 0.00028 $0.0004_{-0.0001}^{+0.0001}$ $0.0003_{-0.0000}^{+0.0000}$
$\varphi_{c}$ [rad] 0.5412 $3.1765_{-1.8449}^{+1.7528}$ $3.6714_{-3.0978}^{+1.6062}$
$d_{L}$ [Mpc] 47054.1 $48698.8670_{-1486.6328}^{+1568.7870}$ $46334.8753_{-891.2652}^{+882.4137}$
$\iota$ [rad] 0.8207 $0.8004_{-0.0207}^{+0.0191}$ $0.8298_{-0.0109}^{+0.0098}$
$\beta$ [rad] 0.0459 $0.1405_{-0.0317}^{+0.0340}$ $0.0401_{-0.0188}^{+0.0188}$
$\lambda$ [rad] 3.3544 $3.3024_{-0.0322}^{+0.0317}$ $3.3350_{-0.0167}^{+0.0188}$
$\psi$ [rad] 1.5634 $1.5680_{-1.5388}^{+0.0442}$ $1.5692_{-1.5613}^{+0.0188}$
A noteworthy phenomenon is that our machine learning pipeline produces a posterior similar to the nested sampling approach in terms of extrinsic parameters, and the posterior distributions of both converge in the same order of magnitude.
In the final part of the results, we explore the computational efficiency of the algorithm. Table 5 shows the speed of generating posterior samples using nested sampling compared to the VPFMPE method. It is important to note that we used nested sampling to search near the true intrinsic parameter values of the injected MBHB. In practice, the time consumed by nested sampling is significantly longer than the times listed in the table. Even under these conditions, our VPFMPE method remains two orders of magnitude faster than the nested sampling method.
Table 5: Comparison of computational time required by different posterior sampling approaches to generate their respective samples.
Number of Posterior Samples Runtime (seconds) Time per Sample
Nested Sampling 11215 3000 0.26750
VPFMPE 14336 240 0.01674
Our model not only prioritizes efficiency but also generates a posterior distribution similar to nested sampling. Moreover, it is well-suited for the majority of MBHB sources anticipated by future Taiji observations. This contribution will significantly enhance global fit analysis.
5 Conclusions
In this study, we have demonstrated the inaugural application of ODE-based generative models in the field of space-based GW detection, taking the parameter estimation of merging MBHBs as a representative example. We recommend the use of ODE-based generative models, particularly the VPFMPE method, as the mainstream machine learning inference approach.
Our method provides comprehensive statistical inference for the 11-dimensional parameters of MBHBs, establishing a new benchmark in this area. This approach, as opposed to likelihood-based inference methods, can lead to both faster training and more accurate estimation.
Furthermore, we propose the integration of model generation strategy based on the symmetry of the system with ODE-based generative models. This enables us to simplify the training task, while still obtaining a model that is applicable to the entire mission period. Note that the coordinate transformation and ODE-based generative models are theoretically applicable to other space-based missions such as LISA.
It is also noteworthy that our method exhibits robustness against challenges such as confusion noise and the complexities of varying reference times, highlighting its reliability and potential for broader applications. The significant speed of our method, surpassing traditional techniques by orders of magnitude, positions it as a valuable tool for preprocessing in global fitting. Moreover, the precise and reliable estimation of sky location parameters (longitude, latitude) proves beneficial in guiding the search for electromagnetic observatories.
This acceleration in data processing and analysis not only streamlines workflows but also broadens the horizons for gravitational wave data exploration. Overall, these advancements herald a transformative era in space-based GW detection, fortifying the capabilities of missions like Taiji and bringing us closer to unraveling the mysteries of the universe.
6 Acknowledgements
This study is supported by the National Key Research and Development Program of China (Grant No. 2021YFC2201901, Grant No. 2021YFC2203004, Grant No. 2020YFC2200100 and Grant No. 2021YFC2201903). International Partnership Program of the Chinese Academy of Sciences, Grant No. 025GJHZ2023106GC.